OpenStudy (anonymous):

find pi or x: sin(pi/6)=(3*sqrt 3)/x

OpenStudy (gorv):

pi =\[\pi\]

OpenStudy (gorv):

sin(pi/6)=1/2

OpenStudy (gorv):

\[\frac{ 1 }{ 2 }=\frac{ 3\sqrt3 }{ x }\]

OpenStudy (gorv):

multiply both side by x ...thn u will get \[\frac{ x }{ 2 }= 3\sqrt 3\]

OpenStudy (gorv):

now multiply boths side by 2